OpenStudy (anonymous):

Can someone tell me wat the answer is to this please as it's been so long out of Skl a3+a4=? And a3-a4=?

OpenStudy (anonymous):

add and subtract like you normally do

OpenStudy (anonymous):

So is that a3+a4=a7 and a3-a4=a-1 then?

OpenStudy (anonymous):

-a1

OpenStudy (anonymous):

for second one

OpenStudy (anonymous):

My teacher said the first one you cannot add and the second you cannot subtract that's why I was so lost

OpenStudy (anonymous):

What exactly is the question?

OpenStudy (anonymous):

To find the answer to both

OpenStudy (anonymous):

All they said was a3+a4?

OpenStudy (anonymous):

Yep so I thought I was wrong to add them both to get a7

OpenStudy (anonymous):

like 3a+4a?

OpenStudy (anonymous):

No a to the power of 3 add to a to the power of 4

OpenStudy (anonymous):

oooh you mean \[a^3+a^4?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[a^3(a+1)\] \[a = -1, 0\]

OpenStudy (anonymous):

Is this what you wanted? The roots?

OpenStudy (anonymous):

You can't add them normally because the exponents are different.

OpenStudy (anonymous):

Wats roots

OpenStudy (anonymous):

a = -1, and 0

OpenStudy (anonymous):

That is your answer for the first one

OpenStudy (anonymous):

\[a^3-a^4 => (1-a)a^3\] \[a = 1, 0\]

OpenStudy (anonymous):

^ Answer for the second one